the sum of two numbers is 156 there product is 900
Found 2 solutions by rajagopalan, venugopalramana:
Answer by rajagopalan(174)

the sum of two numbers is 156 there product is 900
now factorise the product
it composes of 2x2x3x3x5x5 as prime factors
then decompose this into two groups to give a sum 156
for eg 2x3=6 and the balance 2x3x5x5=150 summing 6+150=156
now the answer is 150 and 6.
algebraically also
x*y=900 ..eqn 1
x+y=156...eqn 2
from eqn 1 x=900/y
sub this in eqn 2
(900/y)+y=156 ..eqn3
expanding and solving the quadratic also you get 150 & 6 as solution.

Answer by venugopalramana(3286)

the sum of two numbers(X,Y) is 156 there product is 900
X+Y=156...Y=156-X
XY=900
X(156-X)=900
156X-X^2=900
X^2-156X+900=0
X^2-6X-150X+900=0
X(X-6)-150(X-6)=0
(X-6)(X-150)=0
X=6 AND 150
THE 2 NUMBERS ARE 6 AND 150